OpenStudy (anonymous):

can you apply Rolle's Theorem to the function f(x)=(cos)(pi)(x) on the closed interval [-1/6,1/6]?

OpenStudy (anonymous):

Yes, you can. i.) f(x) is continuous and differentiable on [-1/6, 1/6] ii.) f(-1/6) = f(1/6)

OpenStudy (anonymous):

Therefore there exists some point c such that -1/6 < c < 1/6 and f'(x) = 0

OpenStudy (anonymous):

oops, sorry, and f'(c) = 0

OpenStudy (anonymous):

Thanks!